OpenStudy (anonymous):

Solve for a polynomial with roots -3,6+i

OpenStudy (anonymous):

\[If 6+\iota is aroot then 6-\iota is also aroot.\] required polynomial is \[=\left( x+3 \right)\left( x-6-\iota \right)\left( x-6+\iota \right)\] solve it.

OpenStudy (anonymous):

How do I solve for it?

OpenStudy (anonymous):

\[\left( x-6-\iota \right)\left( x-6+\iota \right)=\left\{ \left( x-6 \right)-\iota \right\}\left\{ \left( x-6 \right)+\iota \right\}\] \[=\left( x-6 \right)^{2}-\iota ^{2}=x ^{2}-12x+36+1=x ^{2}-12x+37\] \[\left( x+3 \right)\left( x-6-\iota \right)\left( x-6+\iota \right)=\left(x+3 \right)\left( x ^{2}-12x+37 \right)\] \[=x ^{3}-12x ^{2}+37x+3x ^{2}-36x+111=x ^{3}-9x ^{2}+x+111\]